OpenStudy (anonymous):

A speed of train A is 20 mph slower than the speed of train B. Train A travels 230 miles in the same time it take train B to travel 330 miles. Find the speed of each train. A is mph B is mph

OpenStudy (anonymous):

A is 40 mph B is 60 mph

OpenStudy (anonymous):

ohh wait no nvm hold on

OpenStudy (anonymous):

idk haha

OpenStudy (phi):

call speed of train B, b and speed of train A, a a= b-20 also speed*time= distance, or time= distance/speed in the same time we have distance for train A and B, and speed: 230/(b-20) = 330/b solve for b